The stone block has a mass of 592.6g and a specific heat capacity of 1200J/Kg°C. A heater supplies 200W of heat energy for 5 minutes. What would the increase in temperature be?
1
Expert's answer
2020-11-10T09:57:06-0500

The quantity of heat needed to produce the increase in temperature of the stone block can be found from the formula:


"Q=mc\Delta T,"

here, "Q" is the quantity of heat needed to produce the increase in temperature, "\Delta T", of the stone block; "m=0.5926\ kg" is the mass of the stone block; "c=1200\ \dfrac{J}{kg\cdot\! ^{\circ}C}" is the specific heat capacity of the stone block.

The quantity of heat supplied by a heater in time "t" can be found from the formula:


"Q=Pt,"

here, "P" is the power of the heater.

Finally, we can equate these two expressions and calculate "\Delta T":


"Pt=mc\Delta T,""\Delta T=\dfrac{Pt}{mc},""\Delta T=\dfrac{200\ W\cdot 5\ min \dfrac{60\ s}{1\ min}}{0.5926\ kg\cdot 1200\ \dfrac{J}{kg\cdot\! ^{\circ}C}}=84.4\ ^{\circ}C."

Answer:

"\Delta T=84.4\ ^{\circ}C."